OpenStudy (anonymous):

how do i simplify this? 4(2x+5)+7

OpenStudy (aroub):

by distributing first..do you know how?

OpenStudy (anonymous):

no, i need help on how to :(

OpenStudy (alfie):

a(b+c) = ab + ac... 4(2x+5) + 7 = 4(2x) + 4(5) + 7

OpenStudy (anonymous):

is that my final answer? or do i simplify it more?

OpenStudy (anonymous):

8x+27?!

OpenStudy (anonymous):

Correct!!